OpenStudy (anonymous):

Given: 2x – 3(x – 5 + x) = 2x – 3 Prove: x = 3

OpenStudy (anonymous):

Mathematical Statement Justification

OpenStudy (anonymous):

2x – 3(x - 5 + x) = 2x - 3 Given 2x – 3(x + x – 5) = 2x – 3 Commutative Property of Addition 2x – 3(2x – 5) = 2x – 3 Combine Like Terms 2x – 6x + 15 = 2x – 3 ? (2x – 6x) + 15 = 2x – 3 ? – 4x + 15 = 2x – 3 ? – 6x + 15 = – 3 Subtraction Property of Equality – 6x = – 18 Subtraction Property of Equality x = 3 Division Property of Equality What are the missing justification steps, in order?

OpenStudy (anonymous):

Distributive Property, Associative Property of Addition, Combine Like Terms Subtraction Property of Equality, Associative Property of Addition, Addition Property of Equality Subtraction Property of Equality, Commutative Property of Addition, Addition Property of Equality Distributive Property, Commutative Property of Addition, Combine Like Terms

OpenStudy (anonymous):

i got Distributive Property, Associative Property of Addition, Combine Like Terms?